An antiproton is a proton antiparticle. The antiparticle of a particle has the same mass as the particle, but with the opposite charge. When an antiproton encounters a proton, annihilation occurs: the particles disappear and gamma radiation is generated. Therefore, when studying antiprotons, it must be possible to keep the antiprotons separate from ordinary material. This is accomplished by putting the accelerated antiprotons in a magnetic field to orbit a very good vacuum. How large must the diameter of such a vacuum chamber be in order to store antiprotons at a speed of 16 km / s in a magnetic field? The magnetic flux density of the magnetic field is 1.9 mT. 




1
Expert's answer
2021-08-31T08:58:55-0400

"F=qvB=4.9\cdot10^{-18}~N,"

"F=mg=\rho Vg=\frac{\pi d^3}6\rho g,"

"d=\sqrt[3]{\frac{6F}{\pi \rho g}}=2900~m."